Summary table Sectors-Plane at 28.06.2011 (files collected at 18.00h)

  Plane I Plane II Plane III Plane IV
Sectors 1 - OK    
Sectors 2 - OK OK* OK*
Sectors 3 - OK    
Sectors 4 - to check 1 trigger cable OK* OK*
Sectors 5 - OK   1 MB missing
Sectors 6 - OK   1 MB missing

OK* = checked before mounted on the frame. TO be re-checked

Logbook of the OEPB and MB which had a failure and had to be exchanged during the commissioning

OEPB logbook

OEPB Address OEPB Nr Exchanged with OEPB Comment
2102     probably burn out, no CAL data. Changed OEP      
2134     Problem with CMS cables. Changed CMS cable.      
2141     Done, MB exchanged 0076=>0167      
2143 474 136        
2145 456 498        
2150 145 376 sometimes no trigger after startup. Problem solved      
             
2224 158 148 NO TDC data      
2228 084 111 no noise data and no calibration.      
             
2309 243 114 still no cal data. It started working again (28.06.11, 17.40      
234B 092 279 OEPB broken. If included in the DAQ we have busy since beginning. Exchanging the OEP solved the problem      
2350 143 210 TO DO no cal data      
2353     TO DO, one TDC missing. It started working again (28.06.11, 17.40h)      

Red: not done, in progress;

Green: finished.

-- AttilioTarantola - 14 Jun 2011
Topic revision: r7 - 01 Jul 2011, SergeyYurevich
 
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Hades Wiki? Send feedback
Imprint (in German)
Privacy Policy (in German)